Problem 1. Find the type, transform to normal form, and solve the following PDEs. (1) uxx – 16uyy = 0 - 2uxy + (2) Uxx Uyy =Sample Solution for Problem 1(1): Hyperbolic, wave equation. Characteristic equation y2 – 16 = (y + 4)(y – 4) = 0. New var
Answer #1

Ans of D

Solutibu:- ou is - So, Here, which The gum equation U42 - 61 xy + quyy =0 S = -6, R=1, the coeffecients of uxy, lata respectiHere, So, o Ral, T=9, Sc.6 26d+ 9 ? -3d-34+9=0 (1-3) -311-3) - (4-3) (1-3) = d=3,3 Now, consider + d = o du dy +-3 2-0 = de lNow, AS w = y + 3* JW 3 gy Also, x ou OX 1 ay Now, ду. ду no ox + ou aur By Now, = DW du au + Ju an su 5304 + su 1 leming 9)Hence, au + 9ə*ч sus з ә4 wәU + 3.5*4 + у ә2- ə x2 anaw Now, OIL ди. Биз да + дц , а 27 ә8 ə o8 1) ди ә ( Using au au Әu (оThus, за 2²u да - | | Fu *РЯ the equation 6 given equation 0 9 д% зә Put (10) in the fullsам . + за t +əq дио 2 18 əч цо- ОUthe So, 9 ใน Since, Because of continuity are awa Uaw above equation becomes :- 604 are -18 are bare Su2 Jugaw 2012 дц2 avawHence, due noumal the given equation :- tum of a²u 102 lou o Ans. we Now, Unr. 6 olay 20 Un > 1, will solve the given equatio

For 2

Uax -2uxy tuyg - A Uxx + Buay + Cugy +D 4x +E vytfuza Mx = 1, y = -1 Now choosing another Chereefenste equation comparing witSubstituting these expressions into Augs+ Event Cung + B up + B Un + Fu =Ğ tougn +4 Unn +0.49 +0. un to u=0 = 4 unna Unno u =Solution of first is